OpenStudy (mariebrittany07):

A triangle has height 15 inches and area 120 inches .What is the length of its base?

OpenStudy (anonymous):

The formula for a triangle is a = 1/2*(b)*(h). So if we plug that in we get: 120 = 1/2*(b)*(15)... from this we can keep simplifying to get a base of 16 inches:) and to check, you can always try a= 1/2 * (16) * (15) with leaves the area to equal 120 :) hope i helped!